OpenStudy (anonymous):

Simplify v to the -7/11 power.

OpenStudy (saifoo.khan):

\[\huge x^{-\frac{7}{11}} \to \frac{1}{x^{\frac{7}{11}}}\]

OpenStudy (anonymous):

Yup, I've got that. Then I multiply the top and bottom by x to the 4/11th power, right? I got x to the 4/11th power over x as my answer.

OpenStudy (saifoo.khan):

umm..

OpenStudy (saifoo.khan):

My answer is simplified too..